OpenStudy (anonymous):

solve each system of equations 2x+y=5 3x-2y=4

OpenStudy (anonymous):

from 1st equation, 2x+y=5 =>y=5-2x from 2nd equation, 3x-2y=4 =>3x-4=2y =>y=(3x-4)/2 comparing y values, 5-2x=(3x-4)/2 =>2(5-2x)=3x-4 =>10-4x=3x-4 =>10+4=3x+4x =>14=7x =>14/7=x =>x=2 putting the value of x, y=5-2x=5-2(2)=5-4=1 so the answer is x=2, y=1